A gas is at 200K. If we wish to double the rms speed of the molecules of the gas, to what value must we raise its temperature?
1
Expert's answer
2021-02-25T15:04:15-0500

The rms speed of the molecules of the gas can be found as follows:


"v_{rms}=\sqrt{\dfrac{3RT}{M}}."

Let the initial rms speed of the molecules of the gas will be:


"v_{rms,1}=\sqrt{\dfrac{3RT_1}{M}}."

Let the final rms speed of the molecules of the gas will be:


"v_{rms,2}=\sqrt{\dfrac{3RT_2}{M}}"

Then, dividing the first expression by the second one and assuming that "v_{rms,2}=2v_{rms,1}", we get:


"\dfrac{v_{rms,1}}{2v_{rms,1}}=\sqrt{\dfrac{T_1}{T_2}},""\dfrac{1}{4}=\dfrac{T_1}{T_2},""T_2=4T_1=4\cdot200\ K=800\ K."